Marigold Creations sells window treatments (shades, blinds, and awnings) to both commercial and residential customers. The following information relates to its budgeted operations for the current year.

Commercial

Residential

Revenues $356,500 $536,000
Direct materials costs $35,000 $50,000
Direct labor costs 140,000 260,000
Overhead costs 91,500 266,500 251,000 561,000
Operating income (loss) $90,000 $(25,000)

The controller, Peggy Kingman, is concerned about the residential product line. She cannot understand why this line is not more profitable given that the installations of window coverings are less complex for residential customers. In addition, the residential client base resides in close proximity to the company office, so travel costs are not as expensive on a per client visit for residential customers. As a result, she has decided to take a closer look at the overhead costs assigned to the two product lines to determine whether a more accurate product costing model can be developed. Here are the three activity cost pools and related information she developed:

Activity Cost Pools

Estimated Overhead

Cost Drivers

Scheduling and travel $91,500 Hours of travel
Setup time 91,000 Number of setups
Supervision 160,000 Direct labor cost
Estimated Use of Cost Drivers per Product

Commercial

Residential

Scheduling and travel 800 700
Setup time 450 250
Compute the activity-based overhead rates for each of the three cost pools. (Round overhead rate for supervision to 2 decimal places, e.g. 0.38.)

Overhead Rates

Scheduling and travel $

  

per hour
Setup time $

per setup
Supervision $

per dollar
Determine the overhead cost assigned to each product line.

Commercial

Residential

Scheduling and travel $

$

Setup time $

$

Supervision $

$

Total cost assigned $

$

Compute the operating income for each product line, using the activity-based overhead rates.

Operating income (loss)

Commercial $

Residential $
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Marigold Creations Answer a
Calculation of ABC rates for overhead A B C D=B+C E=D/A
Activity cost pool Cost Driver Cost Commercial Residential Total Cost Driver Units ABC rates
Scheduling and travel Hours of travel      91,500.00          800.00            700.00            1,500.00              61.00
Setup time Number of setups      91,000.00          450.00            250.00               700.00            130.00
Supervision Direct labor cost 160,000.00 140,000.00    260,000.00       400,000.00                0.40
Total Cost assigned 342,500.00
See E See B F=B*E See C G=C*E
Allocation table Commercial Residential
Activity cost pool Cost Driver ABC Rates Activity used Cost assigned Activity used Cost assigned
Scheduling and travel Hours of travel             61.00          800.00      48,800.00               700.00      42,700.00
Setup time Number of setups           130.00          450.00      58,500.00               250.00      32,500.00
Supervision Direct labor cost                0.40 140,000.00      56,000.00       260,000.00    104,000.00
Total Cost assigned    163,300.00    179,200.00
Operating income Statement Commercial Residential
Revenues                              356,500.00 536,000.00
Less:
Direct Materials                                35,000.00      50,000.00
Direct Labor                              140,000.00 260,000.00
Overhead Costs                              163,300.00 179,200.00 See H
Operating income                                18,200.00     46,800.00